Please help me find a third-degree polynomial equation with rational coefficients that has roots -4 and 2+i.
Answer by stanbon(75887)

Please help me find a third-degree polynomial equation with rational coefficients that has roots -4 and 2+i.
-------------
Roots are -4, 2+i, 2-i
-----
f(x) = (x+4)((x-2)-i)((x-2)+i)
=====
f(x) = (x+4)((x-2)^2 +1)
----
f(x) = (x+4)(x^2 - 4x +5)
------------
f(x) = x^3 -4x^2 + 4x^2 + 5x - 16x + 20
-------
f(x) = x^2 - 11x + 20
---------------
Cheers,
Stan H.
----------